OpenStudy (anonymous):

I need the answer fast! Five of Ethan's family members went on a 6 and one-fourth hour road trip. If they all shared the drive time equally, how many hours did each family member drive? Answer four-fifths 1 and one-fourth seven-eighths 1 and three-twentieths

OpenStudy (anonymous):

just divide (6+1/4) by (5)

OpenStudy (anonymous):

\(6\tfrac{1}{4}\div 5=\frac{25}{4}\div 5=\frac{25}{4}\times \frac{1}{5}\)

OpenStudy (anonymous):

cancel the 5 with the 25 for your answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{25}{4}\times \frac{1}{5}=\frac{\cancel{25}^5}{4}\times \frac{1}{\cancel 5}=\frac{5}{4}\]